How many mL of chlorine gas can be produced by the electrical decomposition of 15.0 g of sodium chloride at 15oC and 98.0 kPa?


1
Expert's answer
2021-04-21T03:55:07-0400


2Na+Cl2→2NaCl


Molar Mass of NaCl = 58.44

Molar Mass of Cl = 70.91

15/58 = 0.2567

= 0.2567 × 70.91 ÷ 2 = 9.1

9.1 × 15 × 98/20

= 657 ml